We are using distributive property and simplify algebraic expressions but I cant get the answer for this because i do not know if I am right&
& -w+4(w+3)
1
Expert's answer
2011-08-30T05:24:33-0400
-w+4(w+3) = -w + 4w + 4*3 = 3w + 12